OpenStudy (anonymous):

A box contains seven action, six comedies, eight dramas, and five science fiction films. Find the probability of choosing a comedy, putting it back, a science fiction, putting it back and then choosing a comedy. The answer choices are: 45/4394 or 30/2197

OpenStudy (jack1):

7+6+8+5 = 26 movies P(comedy then sfiction then comedy) = P (comedy) x P (sci fi) x P (comedy) \(\large P = \frac {7}{26} \times \frac {5}{26} \times \frac {7}{26}\) =???